OpenStudy (butterflyhope):

A Laundromat owner bought eleven new washing machines and eight new dryers. He paid the same amount for a dryer as for a washer. His final cost was $10,450. What was the cost per machine? Thank you :)

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

Im not too sure

OpenStudy (anonymous):

Sorry, cant help you

OpenStudy (butterflyhope):

It's fine :P

OpenStudy (anonymous):

dont give up

OpenStudy (anonymous):

surely someone else will join this room

OpenStudy (butterflyhope):

Kk :)

OpenStudy (anonymous):

Let w equal the cost of a washing machine. Since there are 11 washing machines we have this: 11w +cost of the dryers = total cost of both machines We are given that the cost of dryer is the same as the cost of a washing machine, and eight dryers where purchased. So the cost of the dryers is 8w So our equation looks like this: 11w + 8w = total cost of the both machines We are given that the total cost of both machines is $10,450 So our equation looks like this: 11w + 8w = 10,450 Solve for w. 19w=10,450 w=10,450/19 w=550 So the cost per machine is $550